OpenStudy (anonymous):

Solve by completing the square. .05x^2-x+1=y

OpenStudy (anonymous):

No. There is no completing the square method there.

OpenStudy (anonymous):

steps (0.05 x-1.) x+1 = y (0.223607 x-4.23607) (0.223607 x-0.236068) = y 0.05 x^2-x-y+1 = 0

OpenStudy (anonymous):

u can ask someone else for me also not getting

OpenStudy (anonymous):

Okay.

OpenStudy (anonymous):

satellite73 is best in maths u can call him

OpenStudy (anonymous):

Okay, thanks.

OpenStudy (lgbasallote):

divide everything by 0.5 first

OpenStudy (anonymous):

you can complete the square, first multiply by 2

OpenStudy (lgbasallote):

the coefficient of x^2 needs to be 1 only so you need to divide by 0.5 or in other words...multiply by 2 :|

OpenStudy (anonymous):

in order to complete the square, the first step is to make sure the leading coefficient is 1

OpenStudy (anonymous):

what @lgbasallote is saying

OpenStudy (anonymous):

on the other hand, now that i look at the question, it reads \[.05x^2-x+1=y\] not \[.05x^2-x+1=0\] so it is not clear what "solving" means in this context

OpenStudy (anonymous):

and on further reading, i see that it is \(.05\) and not \(.5\) so you would need to multiply by 20 not 2

OpenStudy (anonymous):

that is if you wanted so solve \[.05x^2-x+1=0\] for \(x\) by completing the square, the first step would be multiply by 20 to make the leading coefficient 1 and write \[x^2-20x+20=0\] then \[x^2-20x=-20\] \[(x-10)^2=-20+100=80\] \[x-10=\pm\sqrt{80}=\pm4\sqrt{5}\] \[x=10\pm4\sqrt{5}\]

OpenStudy (anonymous):

x= any - number, y=0.05x^2-x+1